write the standard form of the equation through the points (4,2) and perpendicular to y=2x-1
Answer by Maths68(1474)

Standard Form of Equation of the line:
y=mx+b
Given
y=2x-1
Compare above equation with the standard form equation
m=2 and b=-1
Since lines are perpendicular multiplicatin of their slope will be (-1)
So slope of the required line will be (-1/2)
Now we have a point(4,2) and slope (-1/2)of the line we can easily find required lines by putting these values in the equation of the straight line poin-slope form.
m=(y2-y1)/(x2-x1)
-1/2=(y-2))/(x-4)
-(x-4)=2(y-2)
-x+4=2y-4
-2y=x-4-4
-2y=x-8
-2y/-2=(x-8)/-2
y=-x/2+8/2
y=-x/2+4
Above equation is the standard form of the equation of the line